Katie bicycled uphill into the mountains at 12 miles per hour and then returned downhill at 20 miles per hour. Her return trip took 2 hours less than er trip up. How long did Katie bicycle uphill? How far was the trip?
Answer by mananth(16946)

distance = x
Time downhill x/ 20
Time uphill x/ 12
extra time while uphill = 2.00 hour
x/ 12 - x/ 20 = 2.00

LCD = 7350 240
multiply by LCD
20 - 12 x = 2.00 * 240
8 x= 480
/ 8
x= 60 niles
Time uphill = 60/12=5 hours